OpenStudy (anonymous):

the equation of the line : 9x-2y=7 (7,-1)?

OpenStudy (anonymous):

9(7)-2(-1)=7 now solve

OpenStudy (anonymous):

I did that and I got 61.

OpenStudy (anonymous):

yeah. scratch what I said. you first need to find the slope of the equation (set 9x-2y=7 equal to y) and then do some other stuff........look at this link for an example: http://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.226304.html

OpenStudy (anonymous):

Im stuck at y-(-9)=(8/9)(x-8)